OpenStudy (anonymous):

find the distance between the points (4,3) and (11,4)

OpenStudy (anonymous):

@BweadedChicken

pooja195 (pooja195):

\[\huge d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\] \[\huge d=\sqrt{(11-4)^2+(4-3)^2}\]

OpenStudy (anonymous):

sorry didnt notice, i was too busy chatting haha, but he showed u the formula Drkid

OpenStudy (anonymous):

SHE****

OpenStudy (anonymous):

so now?

OpenStudy (anonymous):

cant say, pooja wants to ban me

OpenStudy (anonymous):

just help me lol

pooja195 (pooja195):

Im not a mod.

pooja195 (pooja195):

Im just trying to keep you out of trouble @drkid15 if you are here for answers you have come to the wrong site

OpenStudy (anonymous):

Then why did u say (3rd time) that makes me feel like, ur gonna ban me xD

pooja195 (pooja195):

Coz thats the 3rd time i've told you.

OpenStudy (anonymous):

well, fine just because i like u pooja, i wont directly give answers

OpenStudy (anonymous):

ok help me now plz

OpenStudy (anonymous):

\[D = \sqrt{(11 - 4)^{2} (4-3)^{2}}\]

OpenStudy (anonymous):

How do you get rid of those exponents?

OpenStudy (anonymous):

50?

OpenStudy (anonymous):

No, a square root ontop of an exponent throws away exponents, so now its just D = (11 - 4) + (4 - 3)

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

Need any help further?

OpenStudy (anonymous):

so 8

OpenStudy (anonymous):

Yeah